I've tested the Trigger Supervisor with a cosmics trigger.  CODA does
indeed see data.  I didn't actually look at ADC/TDC spectra, but nothing
should have changed there.
        The TS is set up in software to be in non-strobed mode, so that we can
do timing studies.  However, I have found that a cable still needs to be
connected to the CS input of the TS.  Moreover, the cable needs to be
upside down (i.e., the yellow pin is the + side of the ECL, but it needs
to go to the - side of the TS CS input).  I didn't see any pulse in the
CS cable, just a DC offset.  The timing of the Level 1 Accept seems to
be given by the physics trigger.
        There are instructions in the logbook to go back to a common-strobed
trigger.
                                        Karen
Karen Dow wrote:
> 
>         I've replaced the Trigger Supervisor and the readout controller for
> it.  Everything seems to work, although I haven't tried actual triggers,
> as CCR was tuning the beam.  But at least the TS now shows the green GO
> light when I press RunControl GO.
> 
>         I'll try triggers on cosmics in a bit, but I think we can assume shifts
> are back on.
> 
>                                         Karen
This archive was generated by hypermail 2.1.2 : Mon Feb 24 2014 - 14:07:28 EST